Let X be a random variable with probability distribution

X -1 0 1 2 3

F(X) 0.125 .50 0.20 0.05 0.125

a) Find E(X) and VAR(X).

b) Find the probability distribution of the random variable Y= 2X+1. Using the 

probability distribution of Y determine E(Y) and VAR(Y).


1
Expert's answer
2022-06-02T14:58:23-0400

a)


"E(X)=0.125(-1)+0.50(0)+0.20(1)"

"+0.05(2)+0.125(3)=0.55"

"E(X^2)=0.125(-1)^2+0.50(0)^2+0.20(1)^2"

"+0.05(2)^2+0.125(3)^2=1.65"

"Var(X)=E(X^2)-(E(X))^2"

"=1.65-0.55^2=1.3475"

b)


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 y & -1 & 1 & 3 & 5 & 7 \\ \hline
 F(Y) & 0.125 & 0.50 & 0.20 & 0.05 & 0.125 \\
\end{array}"


"E(Y)=0.125(-1)+0.50(1)+0.20(3)"

"+0.05(5)+0.125(7)=2.1"

"E(Y)=2E(X)+1"

"E(Y^2)=0.125(-1)^2+0.50(1)^2+0.20(3)^2"

"+0.05(5)^2+0.125(7)^2=9.8"

"Var(Y)=E(Y^2)-(E(Y))^2"

"=9.8-2.1^2=5.39"

"Var(Y)=(2)^2Var(X)"